Question 1:
You are thinking of investing in Ski Sports, Inc. You have only the following information on the firm at year-end 2018: net income= $50,000, total debt = $1 million, and debt ratio= 70 percent. What is ski's ROE for 2018?
Answer #1

Debt ratio = Debt/assets

0.7 = 1/assets

assets = 1.42857

Equity = assets-debt = 1.42857-1= 0.42857

ROE = net income/equity = 0.05/0.42857= 11.67%